OpenStudy (anonymous):

Write the composite function in the form f(g(x)). [Identify the inner function u = g(x) and the outer function y = f(u).] y = e^9sqrtx I'm lost with this :/

OpenStudy (anonymous):

e^x is f(x), and 9sqrtx is g(x)

OpenStudy (anonymous):

I got that part.

OpenStudy (anonymous):

just dont know what to do after :(

OpenStudy (anonymous):

uhhhh...

OpenStudy (anonymous):

multiply them?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

set 9sqrtx=u i think

OpenStudy (anonymous):

yes. y=f(u)=e^u & u=g(x)=9sqrtx

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

let me input it!

OpenStudy (anonymous):

what is the question asking?

OpenStudy (anonymous):

did you get the screen shot I took?

OpenStudy (anonymous):

I got the first part.

OpenStudy (anonymous):

ok, so take the derivative now using chain rule

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (anonymous):

Yes, let me look at my notes. We just started doing that.

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

college. Cal 1

OpenStudy (anonymous):

I'm a freshman in high school

OpenStudy (anonymous):

how is calc? is it hard?

OpenStudy (anonymous):

wow. That's weird. Yes it is a little difficult since my teacher has a thick accent and his teaching methods arent clear

OpenStudy (anonymous):

:) anyways back to the question, you take the derivative of e^9sqrtx normally, then multiply by the derivative of 9sqrtx

OpenStudy (anonymous):

so it would be e^9sqrtx(9sqrtx)?

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

but it's not multiplied by normal 9sqrtx, it's multiplied by the derivative of 9sqrt

OpenStudy (anonymous):

9sqrtx*

OpenStudy (anonymous):

and the derivative of 9sqrtx is 1/2sqrtx

OpenStudy (anonymous):

so e^9sqrtx(9(1/2sqrtx))?

OpenStudy (anonymous):